OpenStudy (anonymous):

Find the normal line to the curve x^2 + 2y = 8 andperpendicular to the line x -2y + 4 = 0

OpenStudy (amistre64):

you need to determine the derivative

OpenStudy (amistre64):

since a normal is perpendicular to a tangent, and derivatives define tangents .... find where the derivative is equal to the slope of the given line

OpenStudy (anonymous):

can you please just say the steps on how to solve it and i'll be the one to solve it...

OpenStudy (amistre64):

those are the steps

OpenStudy (anonymous):

that only?

OpenStudy (amistre64):

yep .... determine when the derivative is equal to the slope of the line; this finds the point that you would need to use in a point slope form of the normal line equation

OpenStudy (anonymous):

i think i might get a hard time with that steps coz i'm new to this topic.. can you please elaborate further?

OpenStudy (amistre64):

tell me what the slope of the given line is ...

OpenStudy (anonymous):

you can all help me :)

terenzreignz (terenzreignz):

^As soon as you're done with that. (What's the slope of the line given?)

OpenStudy (anonymous):

1/2

OpenStudy (amistre64):

good, then we will need to equate that to the derivative, and also, the normal line will perp that slope. what is the perpendicular slope of 1/2 ? for future use

OpenStudy (anonymous):

-2

OpenStudy (amistre64):

youre doing great; lets form the normal line equation up so far: (y-a) = -2(x-b) ... just need to define the point (a,b) so lets move on to the derivative what do you get as the derivaitve of the given curve?

OpenStudy (amistre64):

the derivative of: x^2 + 2y = 8 give it your best attempt

OpenStudy (anonymous):

2x + 2 y1 = 0 dy/dx = -x

OpenStudy (amistre64):

good job ... so we need to determine when dy/dx = 1/2 -x = 1/2 when x=?

OpenStudy (anonymous):

-1/2

OpenStudy (amistre64):

pretty simple eh :) so we have part of our point (a,b) ... (-1/2, b) lets find b (the y value) by plugging in x=-1/2 into the curve (-1/2)^2 + 2y = 8 , solve for y

OpenStudy (anonymous):

y = 31/8

OpenStudy (amistre64):

marvolous ... so (a,b) = (-1/2, 31/8)

OpenStudy (amistre64):

can you finish it from there?

OpenStudy (anonymous):

i just need to use point slope form right?

OpenStudy (amistre64):

im not sure what format they want you to present it in, but yes ... the point slope form of a line is a line equation.

OpenStudy (anonymous):

i'm just finding the equation so that may be good.. thanks for the help! :)

OpenStudy (amistre64):

youre welcome ;)